OpenStudy (anonymous):

The protons of an atom are A. negatively charged and located inside the nucleus. B. positively charged and located inside the nucleus. C. negatively charged and located outside the nucleus. D. positively charged and located outside the nucleus.

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

easy way to remember protons- sounds positive electrons- negative neutrons-neutral Found in nucleus are protons & neutrons. On the outer shells is the electrons.

OpenStudy (anonymous):

Thank you very much! That is an easy way to remember.

OpenStudy (anonymous):

no problem, glad I helped u.

OpenStudy (anonymous):

Could you help me with a few more questions?

OpenStudy (anonymous):

sure, if I'll try to but be quick because I have to go to sleep real soon.

OpenStudy (anonymous):

I'll try to*

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Describe how protons, neutrons, and electrons each affect the charge and mass of an atom.

OpenStudy (anonymous):

well protons are positively charged, neutrons neutral(no charge) and electrons negatively charged. So, charges are affected by protons and electrons. Also, mass is affected by all.

OpenStudy (anonymous):

Thank you for explaining that. :)

OpenStudy (anonymous):

no problem.

OpenStudy (anonymous):

You have time for one more? An atom of boron has an atomic number of 5 and an atomic mass of 11. The atom contains A. 5 protons, 6 electrons, and 5 neutrons. B. 11 protons, 11 electrons, and 5 neutrons. C. 5 protons, 5 electrons, and 6 neutrons. D. 11 protons, 6 electrons, and 6 neutrons.

OpenStudy (anonymous):

all of them have different masses, and electrons have the smallest mass, some people consider it to be 0, since it's so small.

OpenStudy (anonymous):

So what's the answer then?

OpenStudy (anonymous):

1. An element's or isotope's atomic number tells how many protons are in its atoms. 2. An element's or isotope's mass number tells how many protons and neutrons in its atoms. no. of neutrons + no. of protons = Mass number Atomic number 5 tells us it has 5 protons x + 5 = 11 x = 11- 5 x = 6 so 5 protons, 6 neutrons and 5 electrons. C.

OpenStudy (anonymous):

answer is C, and remember for now no. of protons will be the same as no.of neutrons.

OpenStudy (anonymous):

Oh okay! Thank you! :)

OpenStudy (anonymous):

and atomic mass is total of protons + neutrons.

OpenStudy (anonymous):

Oh okay.

OpenStudy (anonymous):

ok, I got to go now Kayla.

OpenStudy (anonymous):

Thank you so much for your help!

OpenStudy (anonymous):

just try your best, and google it if you have to.

OpenStudy (anonymous):

no problem.

OpenStudy (anonymous):

I will. Thank you. Have a good day!

OpenStudy (anonymous):

well it's goodnight for me over here lol, but yeh, u have a good day.

OpenStudy (anonymous):

Oh! Good night! lol.